OpenStudy (lizz123):

Is this correct: y=6x-33 Write an equation in slope intercept form the line passing through each pair of points. (6,3)(-6,9)

OpenStudy (haleyelizabeth2017):

Give me a second to see :)

OpenStudy (lizz123):

k

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 6}}\quad ,&{\color{blue}{ 3}})\quad &({\color{red}{ -6}}\quad ,&{\color{blue}{ 9}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (haleyelizabeth2017):

She is wanting slope-intercept form lol

OpenStudy (haleyelizabeth2017):

It is incorrect. Try again using the slope formula @jdoe0001 gave you ;)

OpenStudy (lizz123):

\[m=\frac{ 9-3 }{ -6-6} =\frac{ 6 }{ 0 }=6\]

OpenStudy (haleyelizabeth2017):

-6-6 is not 0.

OpenStudy (haleyelizabeth2017):

And even if it was, that would not work....anything over 0 does not work, but 0 over anything does.

OpenStudy (lizz123):

-6-6= -12

OpenStudy (haleyelizabeth2017):

correct!

OpenStudy (haleyelizabeth2017):

so 6/-12 simplifies to what?

OpenStudy (lizz123):

1/2

OpenStudy (lizz123):

simplified

OpenStudy (haleyelizabeth2017):

close....-1/2

OpenStudy (lizz123):

ohh now \[y=-\frac{ 1 }{ 2 }x+b\]

OpenStudy (haleyelizabeth2017):

yes. now plug in one ordered pair, and what do you get for b?

OpenStudy (lizz123):

(6,3)

OpenStudy (lizz123):

\[3=-\frac{ 1 }{ 2 }(6)+b\]

OpenStudy (haleyelizabeth2017):

correct. now what is -1/2*6?

OpenStudy (lizz123):

\[3=\frac{ 6 }{ 12 }\]

OpenStudy (haleyelizabeth2017):

6/12? How'd you get that?

OpenStudy (lizz123):

I always get confused on this part. I multiplied

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 6}}\quad ,&{\color{blue}{ 3}})\quad &({\color{red}{ -6}}\quad ,&{\color{blue}{ 9}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ 9}}-{\color{blue}{ 6}}}{{\color{red}{ -6}}-{\color{red}{ 6}}}\implies \cfrac{\cancel{ 6}}{\cancel{ -12}}\implies -\cfrac{1}{2} \\ \quad \\ % point-slope intercept y-{\color{blue}{ 3}}={\color{green}{ -\cfrac{1}{2}}}(x-{\color{red}{ 6}})\implies y-3= -\cfrac{1}{2}x+\left(-\cfrac{1}{2}\cdot -6\right) \\ \quad \\ y-3= -\cfrac{1}{2}x+\left(-\cfrac{-6}{2}\cdot \right)\implies y-3= -\cfrac{1}{2}x+\left(\cfrac{\cancel{ -6}}{\cancel{ -2}} \right)\)

OpenStudy (haleyelizabeth2017):

lol -(6/2) what is that?

OpenStudy (haleyelizabeth2017):

-(6/2) is just another way of saying -1/2(6)

OpenStudy (haleyelizabeth2017):

@jdoe0001, would you mind finishing using the slope-intercept form so it's easier for her?

OpenStudy (jdoe0001):

anyhow -6/-2 = +3 thus \(\bf slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ 9}}-{\color{blue}{ 6}}}{{\color{red}{ -6}}-{\color{red}{ 6}}}\implies \cfrac{\cancel{ 6}}{\cancel{ -12}}\implies -\cfrac{1}{2} \\ \quad \\ % point-slope intercept y-{\color{blue}{ 3}}={\color{green}{ -\cfrac{1}{2}}}(x-{\color{red}{ 6}})\implies y-3= -\cfrac{1}{2}x+\left(-\cfrac{1}{2}\cdot -6\right) \\ \quad \\ y-3= -\cfrac{1}{2}x+\left(-\cfrac{-6}{2}\cdot \right)\implies y-3= -\cfrac{1}{2}x+\left(\cfrac{\cancel{ -6}}{\cancel{ -2}} \right) \\ \quad \\ y-3= -\cfrac{1}{2}x+3\implies y\cancel{ -3+3}=-\cfrac{1}{2}x+3+3 \\ \quad \\ y=-\cfrac{1}{2}x+6\)

OpenStudy (lizz123):

Thanks @haleyelizabeth2017 and @jdoe0001 I now understand!!

OpenStudy (jdoe0001):

yw

OpenStudy (haleyelizabeth2017):

You are welcome.... :) but @jdoe0001 did most of the work :)

OpenStudy (lizz123):

:D